OpenStudy (bacon):

Please Help! (3w+5)^2

OpenStudy (ranga):

(x+y)^2 = x^2 + 2xy + y^2

OpenStudy (bacon):

so would that be the FOIL method?

OpenStudy (ranga):

The formula in my first reply is an algebraic identity. You need the foil method only when you want to multiply out (3w+5)(3w+5). Either way you will get the same answer after simplification. But the identity method will be faster.

OpenStudy (bacon):

if it wont take too much time, can you tell me what the identity method is?

OpenStudy (ranga):

(x+y)^2 = x^2 + 2xy + y^2 The identity says: square the first term + 2 * (first term) * (second term) + square the second term. (3w + 5)^2 = (3w)^2 + 2 * 3w * 5 + (5)^2 = ?

OpenStudy (bacon):

ok thanks

OpenStudy (bacon):

so i got 9w+30w+25

OpenStudy (ranga):

(3w + 5)^2 = (3w)^2 + 2 * 3w * 5 + (5)^2 = 9w^2 + 30w + 25

OpenStudy (bacon):

ok thanks so much :D

OpenStudy (ranga):

You are welcome. If you FOIL it and simplify you will get the same answer: (3w+5)^2 = (3w+5) * (3w+5) = (3w)(3w) + (3w)(5) + (5)(3w) + (5)(5) = 9w^2 + 15w + 15w + 25 = 9w^2 + 30w + 25.